OpenStudy (anonymous):

how do you factor 5x^2-11x+36=0 and 3x^2-15x-18?

OpenStudy (anonymous):

are you sure you typed the 1st equation right? On 2nd factor out 3

OpenStudy (anonymous):

um.. how?

OpenStudy (anonymous):

3x^2-15x-18=3*x^2-3*5x-3*6=3(x^2-5x-6)

OpenStudy (anonymous):

I don't understand how you can find x.

OpenStudy (anonymous):

3x^2-15x-18=3(x^2-5x-6)=3(x-6)(x+1) x=6 or x=-1

OpenStudy (anonymous):

Here you could find good explanation http://www.purplemath.com/modules/factquad.htm